OpenStudy (anonymous):

Use the formula A = lw to find the width of a rectangle whose area is 46.875 square meters and whose length is 7.5 meters. Round to two decimal places. Enter the numeric answer only.

OpenStudy (campbell_st):

substitute the known information into your given formula \[46.875 = 7.5 \times w\] just solve for w.

OpenStudy (anonymous):

A= length * hight A=46.875 m^2 l=7.5 m w= unknown so 46.875=7.5*w w=46.875/7.5